Allison:

Simplifying Rational Exponents 2.0

Allison:

(-32)^3/5

Allison:

-8?

justjm:

It would be converted to \( \sqrt[5]{\left(-32\right)^{3}}\) You could that by first evaluating \( (-32)^{\frac{1}{5}} \) which is -2 and then cube it so it's -8

justjm:

yup!

Allison:

Thank youuu :))

justjm:

My old Algebra 1 teacher from several years ago taught me a nice remembering cue for this. It looks pretty stupid but I don't think I've done any problems like this without her remembering cue. It goes like this "Power over root, power over root, power over root, root, root, root, roooooot" if that would help lol

Allison:

Hahah that’s so cute, thank you for sharing :)